40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 734     Entry time: Thu Jul 24 11:49:07 2008
Author: Masha 
Type: Summary 
Category: Auxiliary locking 
Subject: belated weekly summary 
I designed a high pass filter to whiten the spectrum from the Mach Zehnder to optimize the
input into the ADC. The swept sine response measurement and the effect of the filter on the
spectrum are attached. If I start using the digital system (it is currently down in Bridge),
I will decide if the filter needs to be improved/better matched to the ADC there.

I moved from the 40m to Rana's lab in bridge. I am making a new and improved Mach Zehnder
setup with a 50m fiber in one arm; currently the transmission through the fiber is 44%. I
am working out how to mode match the laser to the fiber to improve this number.
Attachment 1: filter_tr_function.pdf  4 kB  | Hide | Hide all
filter_tr_function.pdf
Attachment 2: filtered_spectr0724.pdf  9 kB  | Hide | Hide all
filtered_spectr0724.pdf
ELOG V3.1.3-